One channel of my B2 works and adjusts correctly while I cannot adjust the bias current or DC balance on the other channel. The bias just reads 0.000 and won't adjust. The power supply provides proper + / - 85v and I've removed and checked every transistor on the drive board with an Atlas DCA55 checker and all are good. I've also checked all the diodes as well. A visual of all the resistors don't show anything burned out and have checked them in-circuit for open and all seem OK. I also removed the VFET's and checked them per the method described by Ecowars. Have not checked all the caps on the drive board as of yet as I thought I might get some opinions before I went any further.